OpenStudy (dls):

The solution of (x+logy)dy+ydx=0 when y(0)=1 is?

OpenStudy (anonymous):

Looks like an exact equation.

OpenStudy (dls):

Attempt: \[\LARGE (x+\log y)\frac{dy}{dx}=-y\] \[\LARGE \frac{dx}{dy}=-\frac{(x+logy)}{y}\] \[\LARGE \frac{dx}{x+\log y}=- \frac{dy}{y}\] urg doesnt work :/ variable seperable fails.. nor its homogenous.. not sure about linear..

OpenStudy (anonymous):

It's an exact equation.

OpenStudy (anonymous):

There is an \(f(x,y)\) such that:\[ f_xdx+f_ydy=0 \]

OpenStudy (dls):

wait let me try now then

OpenStudy (dls):

thanks for the clue

OpenStudy (anonymous):

why not ask the question in here?

OpenStudy (dls):

xy is there for sure..

OpenStudy (anonymous):

\[ [x\ln(x)]'=\ln(x)+1 \]

OpenStudy (anonymous):

\[ [x\ln(x)-x]'=\ln(x)+1-1=\ln(x) \]

OpenStudy (dls):

\[\LARGE xy+(y \log y-y)\] seems so

OpenStudy (anonymous):

\[+C\]

OpenStudy (dls):

yup

OpenStudy (anonymous):

\[ xy+(y\ln(y)-y)+C=C_2 \]

OpenStudy (dls):

C=0 anyway,got the answer! thanks!